# Routing and wavelength assignment in quantum key distribution networks: power control heuristics for quantum-classical multiplexing

### v1

@@ Abstract
As quantum key distribution networks grow in size and complexity, resource assignment has become increasingly important. In passive optical networks without wavelength conversion, we need to assign a full route between origin and destination with the same wavelength from a finite set. This problem is computationally intensive and the common solution in classical optical networks is using heuristics. We adapt these heuristics to hybrid quantum networks where the quantum channel can share some of the optical links with classical channels. In this quantum-classical multiplexing, nonlinear effects can become the limiting factor in the range of the network. The signal in the classical channels can be subject to Raman Scattering or Four-Wave-Mixing and produce light in the quantum channels. While these effects are not efficient, even a single photon can ruin the quantum channel. We propose heuristics for the routing and wavelength assignment problem for hybrid quantum-classical networks with power control for the classical channels. By keeping the transmitted power to its bare functional minimum, we can reduce the interference to the quantum channels. We study their efficiency under different scenarios.
Index Terms:
Routing and wavelength assignment, QKD networks, Quantum-classical multiplexing.
I Quantum key distribution networks
@@
Communication protocols that use the laws of Quantum Mechanics allow new applications beyond classical possibilities. In particular, the limitations of quantum measurement, where the state of a superposition is destroyed after being observed, can be used to establish an arbitrarly long secret key between two parties that use a private, but insecure, quantum communication channel and a public, but authenticated, classical communication channel. This offers an additional level of security based on the physical principles that can be combined with the usual cryptographic methods based on computational complexity. Quantum Key Distribution, QKD, is a mature quantum technology for which there are multiple protocols, experimental proofs of concept, standards and commercial products [1]. Now that QKD links between two sides can be deployed with relative ease, there is a growing interest in QKD networks with multiple nodes in which any pair of nodes can establish a secret key securely.
@@
As in most quantum communication demonstrations, in order to cover large distances fast, photons are the quantum system of choice. The main proposals are free-space QKD, including satellite communication, and optical fiber QKD networks that take advantage of the existing optical communication network infrastructure. We will concentrate on these networks.
@@
While the first QKD demonstrations used dedicated dark fiber links, the current objective is a gradual integration with existing fiber communication networks. Ideally, QKD networks would be part of flexible Software Defined Networks, where there is a logical control plane deciding which connections can be established and how the network resources are assigned [2].
@@
QKD links need a classical channel for public discussion, including error correction and privacy amplification, and a quantum channel which can basically use most of existing classical optical fiber technology with the important exception of amplifiers. The same principles that prevent an eavesdropper learning a quantum state forbid noiseless amplification and there is a distance limit on QKD. Photons can get lost during transmission and, above a certain loss threshold, we can no longer guarantee the observed imperfections are not due to an adversary that is attacking the system. Current QKD systems, depending on the quality of the detectors or the optical fiber used, can cover distances from tens to hundreds of kilometers. The systems that use standard optical fiber devices are limited to the smaller end of that range.
@@
QKD hardware is chosen to minimize noise and losses. We will concentrate on discrete variable QKD protocols where the signal sent into the channel usually has less than one photon on average. This makes eavesdropping harder as we cannot measure the photon without disturbing its state and we cannot keep a copy for measurement and forward another photon. However, this means that stray photons in the quantum channel band can introduce fatal errors.
@@
Thermal photons are not much of an issue compared to the noise photons that appear in hybrid networks where the same optical fiber carries the classical and the quantum channels needed in QKD or when we share the resources with classical traffic. In these cases, there can be contamination from the classical signal. For wavelength division multiplexing, the main contamination mechanism are nonlinear effects that convert photons from the classical channel bands into the quantum band. In particular, depending on the chosen wavelengths, Raman Scattering and Four-Wave-Mixing can create noise photons from classical channels with moderate powers with a high enough probability to be the limiting factor to the maximum link length [3, 4]. These effects are more important for longer shared paths between the classical and the quantum signals.
@@
II Network description and traffic demands. Routing and wavelength assigment
@@
We consider a series of network demands where we are given an origin and a destination which ask for a lightpath: a path connecting the origin and destination nodes using a dedicated wavelength. Two lightpaths sharing a link must have a different wavelength, but non-overlapping lightpaths can use the same or different wavelengths. These restrictions model what happens in optical networks without wavelength conversion, which is a reasonable scenario for our QKD network.
@@
For a given topology and a series of traffic requests, the routing and wavelength assigment, RWA, problem consists in finding lighpaths for each origin and destination minimizing the number of requests that cannot be fulfilled (for instance if there are no available wavelenghts in any route between the nodes). This problem is known to be NP-complete for its usual formulation [7]. Full optimization is quite involved and, in practice, there is a set of efficient heuristics which approximate the best solution faster and using less comptuational resources [8].
A usual strategy is using the heuristic shortest-path-first-fit, SPFF, and its variations. In the heuristic, we begin using Dijkstra’s algorithm to find the shortest path between origin and destination (or the $k$ shortest paths in many cases). Then, assuming we have an ordered list of available wavelengths, we assign a lightpath to the shortest route which has a free wavelength, assigning the wavelength with the smallest index in our ordering. If no such a path can be found, the connection is blocked (the traffic request is denied).
The RWA and related problems are also relevant to QKD networks and there is an increasing interest in resource allocation methods for different kinds of quantum network [9]. We continue here the line started in [10], which also includes a comparison to other RWA proposals for QKD. The proposed heuristics focus on reducing the problems caused by multiplexing classical and quantum channels. The ultimate purpose is to give methods as simple as possible which can be deployed with little changes in current optical fiber networks.
III Channel model
When routing classical and quantum channels through the same optical fiber, we need to estimate how much noise the classical channels creates in the quantum channels. We will use a simplified model that gives an approximate measure of the expected interference without a complete, and costly, physical simulation that takes into account all the fiber parameters and has accurate models of the relevant nonlinear effects. This way, the heuristics can take into account the limiting noise effects in quantum-classical multiplexing while they are still simple and use few computational resources.
@@ III Channel model
We will use the quantum signal-to-noise ratio, QSNR, in the quantum channel as a guide. In most natural scenarios, a smaller signal-to-noise ratio in the quantum channel is related to a higher QBER and a lower established key rate. We will use a QSNR threshold to determine whether a quantum channel can provide fresh secret bits to a key or not. For each QKD protocol, we can always find out which SNR threshold corresponds to an exchange that is doomed to fail due to a QBER above the minimum value. In general, this is a direct function of the total distance of the quantum link, but, in hybrid QKD networks, the classical part and the noise it induces is also relevant.
III-A Classical and quantum signal to noise ratios
We will start with general definitions of the signal-to-noise ratios we need that will be later refined in the model.
@@ III-A Classical and quantum signal to noise ratios
In the classical channels we use the usual definition for the signal-to-noise ratio
$SNR=\frac{P}{N}$
@@
as the ratio between the signal power $P$ and the noise power $N$. For values below a certain threshold, which depends on the receivers, we assume the signal cannot be interpreted correctly. If we have a fixed power at each transmitter, the receivers at the end of shorter links will receive a higher power. We are going to assume there is a fixed noise power for each link, mostly related to thermal noise at the receivers.
@@
We estimate the effects on the quantum channel by defining a quantum signal-to-noise ratio
$QSNR=\frac{P_{q}}{N}=\frac{P_{q}}{N_{f}+N_{c}}=\frac{P_{o}10^{-\alpha L}}{N_{f}+\sum\gamma_{nl}P_{s}}.$
We consider all the powers are normalized to the single photon level for the corresponding wavelength so that we can interpret them as a photon number. $P_{q}$ gives an estimation of the expected photon number that reaches the receiver in a detection window. It can be computed from the transmitted photon number and the attenuation in the journey to the destination node. We will also include in that term effects that are equivalent to losses, like detection efficiencies below 1, which are common for the simplest single photon detectors. For instance, many systems use avalanche photodiodes which are notoriously inefficient at the IR telecom bands, where 10-20% efficiencies are the norm.
@@
The fixed amount $N_{f}$ that collects thermal photons and equivalent effects, like dark counts in the detectors (false detections caused by phonons or other effects that are not related to the incoming light).
@@
We will give a simplified model to estimate $N_{c}$ as the product of a non-linear coefficient $\gamma_{nl}$ by a sum of the power of the classical power $P_{s}$ of each classical channel that shares a link with the quantum channel normalized by the shared total distance. This is a linearization that does not cover the full physical interaction, but it can give a reasonable estimation of which classical channels are likely to introduce the most noise without a detailed physical model of each situation. The parameters can be adjusted to model each particular scenario, including using different non-linear coefficients for different links using different optical fibers or whether the classical and quantum signals are copropagating or counterpropagating.
IV System Model

### v2

# Lidia Ruiz Juan Carlos Garcia-Escartin

@@ Abstract
As quantum key distribution networks grow in size and complexity, resource assignment has become increasingly important. In passive optical networks without wavelength conversion, we need to assign a full route between origin and destination with the same wavelength from a finite set. This problem is computationally intensive and the common solution in classical optical networks is using heuristics. We adapt these heuristics to hybrid quantum networks where the quantum channel can share some of the optical links with classical channels. In this quantum-classical multiplexing, nonlinear effects can become the limiting factor in the range of the network. The signal in the classical channels can be subject to Raman Scattering or Four-Wave-Mixing and produce light in the quantum channels. While these processes are not efficient, even a single photon can ruin the quantum channel. We propose heuristics for the routing and wavelength assignment problem for hybrid quantum-classical networks with power control for the classical channels. By keeping the transmitted power to its bare functional minimum, we can reduce the interference to the quantum channels. We study their efficiency under different scenarios.
keywords:
QKD , RWA , Hybrid Networks , Power Control
journal: Computer Networks
[…]
[inst1]organization=Departamento de Teoría de la Señal y Comunicaciones e Ingeniería Telemática, Universidad de Valladolid,addressline=Paseo Belén 15, city=Valladolid, postcode=47011, country=Spain
[…]
[inst2]organization=Laboratory for Disruptive Interdisciplinary Science (LaDIS), Universidad de Valladolid,city=Valladolid, postcode=47011, country=Spain
1 Quantum key distribution networks
@@
Communication protocols that use the laws of Quantum Mechanics allow new applications beyond classical possibilities. In particular, the limitations of quantum measurement, where the state of a superposition is destroyed after being observed, can be used to establish an arbitrarily long secret key between two parties that use a private, but insecure, quantum communication channel and a public, but authenticated, classical communication channel. This offers an additional level of security based on physical principles that can be combined with the usual cryptographic methods based on computational complexity. Quantum Key Distribution, QKD, is a mature quantum technology for which there are multiple protocols, experimental proofs of concept, standards and commercial products [1, 2, 3, 4]. Now that QKD links between two sides can be deployed with relative ease, there is a growing interest in QKD networks with multiple nodes in which any pair of nodes can establish a secret key securely.
@@
As in most quantum communication demonstrations, in order to cover large distances fast, photons are the quantum system of choice. The main proposals are free-space QKD, including satellite communication, and optical fiber QKD networks that take advantage of the existing optical communication network infrastructure. We will concentrate on these fiber networks.
@@
While the first QKD demonstrations used dedicated dark fiber links, the current objective is a gradual integration with existing fiber communication networks. Ideally, QKD networks would be part of flexible Software Defined Networks, SDN, where there is a logical control plane deciding which connections can be established and how the network resources are assigned [5, 6, 7].
@@
QKD links need a classical channel for public discussion, including error correction and privacy amplification, and a quantum channel which can basically use most of the existing classical optical fiber technology with the important exception of amplifiers, which must be considered with care. The same principles that prevent an eavesdropper learning a quantum state forbid noiseless amplification and there is a distance limit for QKD links. Photons can get lost during transmission and, above a certain loss threshold, we can no longer guarantee the observed imperfections are not due to an adversary that is attacking the system. Current QKD systems, depending on the quality of the detectors and the optical fiber used, can cover distances from tens to hundreds of kilometers. The systems that use standard optical fiber devices, with simple transmitters and detectors, are limited to the smaller end of that range.
@@
QKD hardware is chosen to minimize noise and losses. We will concentrate on discrete variable QKD protocols where the signal sent into the channel usually has less than one photon on average. This makes eavesdropping harder as we cannot measure the photon without disturbing its state and we cannot keep a copy for measurement and forward another photon. However, this means that stray photons in the quantum channel band can introduce fatal errors. The heuristics we give could also be considered for continuous variable QKD protocols [8] where amplifiers can help in certain situations. We will discuss only discrete variable protocols, which need a higher protection against photon contamination.
@@
Thermal photons are not much of an issue compared to the noise photons that appear in hybrid networks where the same optical fiber carries the classical and the quantum channels needed in QKD or when we share the resources with classical traffic. In these cases, there can be contamination from the classical signal. For wavelength division multiplexing, the main contamination mechanism is the presence of nonlinear effects that convert photons from the classical channel bands into the quantum band. In particular, depending on the chosen wavelengths, Raman Scattering and Four-Wave-Mixing can create noise photons from classical channels with moderate powers with a high enough probability to be the limiting factor to the maximum link length [9, 10, 12, 11, 13]. These processes are more important for longer shared paths between the classical and the quantum signals.
@@
2 Network description and traffic demands. Routing and wavelength assigment
@@
We consider a series of network demands where we are given an origin and a destination which ask for a lightpath: a path connecting the origin and destination nodes using a dedicated wavelength. Two lightpaths sharing a link must have a different wavelength, but non-overlapping lightpaths can use the same or different wavelengths. These restrictions model what happens in optical networks without wavelength conversion, which is a reasonable scenario for our QKD network. Wavelength conversion below the single photon level is technologically challenging and, while it can be considered for the classical signals, it would increase the complexity and cost of the QKD network.
@@
For a given topology and a series of traffic requests, the routing and wavelength assigment, RWA, problem consists in finding lighpaths for each origin and destination, minimizing the number of requests that cannot be fulfilled (for instance if there are no available wavelenghts in any route between the nodes). This problem is known to be NP-complete for its usual formulation [16]. Full optimization is quite involved and, in practice, there is a set of efficient heuristics which approximate the best solution faster and using less computational resources [17].
A widely used strategy is using the heuristic shortest-path-first-fit, SPFF, and its variations. In this heuristic, we begin using Dijkstra’s algorithm to find the shortest path between origin and destination (or the $k$ shortest paths in many cases). Then, assuming we have an ordered list of available wavelengths, we assign a lightpath to the shortest route which has a free wavelength, assigning the wavelength with the smallest index in our ordering. If no such a path can be found, the connection is blocked (the traffic request is denied).
The RWA and related problems are also relevant to QKD networks and there is an increasing interest in resource allocation methods for quantum networks [18], including schemes with trusted or semitrusted relay nodes [19, 20, 21, 22, 18, 23, 24, 25] and with combined time and wavelength multiplexing [26, 27, 28, 29, 19, 20, 21, 25]. We continue here the line started in [30], which also includes a comparison to other RWA proposals for QKD. The proposed heuristics focus on reducing the problems caused by multiplexing classical and quantum channels. The ultimate purpose is to give methods as simple as possible which can be deployed with little changes in current optical fiber networks. The main difference with respect to recent quantum-specific RWA algorithms is that our heuristics explicitly take into account the most limiting physical noise mechanisms in QKD and provide mitigation at the network level. This naturally leads to power control strategies for the classical channels.
3 Link model: nonlinear effects, signal-to-noise ratios and quantum key rate
When routing classical and quantum channels through the same optical fiber, we need to estimate how much noise the classical channels create in the quantum channels. We will use a simplified model for each link that gives an approximate measure of the expected interference instead of a complete, and costly, physical simulation including all the fiber parameters and with accurate models of all the relevant nonlinear effects. This way, the heuristics can approximate the limiting noise effects in quantum-classical multiplexing while they are still simple and use few computational resources.
@@ 3 Link model: nonlinear effects, signal-to-noise ratios and quantum key rate
The effect of the noise will be estimated with the quantum signal-to-noise ratio, QSNR, in the quantum channel. In most natural scenarios, a smaller signal-to-noise ratio in the quantum channel is related to a higher QBER and a lower established key rate. We will use a QSNR threshold to determine whether a quantum channel can provide fresh secret bits to a key or not. For each QKD protocol, we can always find out which QSNR threshold corresponds to an exchange that is doomed to fail due to a QBER above the minimum value. In general, this is a direct function of the total distance of the quantum link, but, in hybrid QKD networks, the noise produced by the classical part is also relevant. We will show the effects on the final key rate with an example for the BB84 protocol under certain restrictions.
3.1 Nonlinear effects
A common solution for quantum-classical multiplexing is keeping the classical channels in the C-band around $1550\text{\,}\mathrm{nm}$ and moving the quantum transmission to the O-band around $1310\text{\,}\mathrm{nm}$. We will work with this channel arrangement, for which the dominant noise source will be Raman scattering [10]. We will be concerned by the tails of the Raman spectrum produced by signals in the C-band. These tails have weak levels in the O-band, but they are noticeable at the quantum level.
Raman scattering occurs when the photons of the classical channel interact with the vibrational modes of the optical fiber. There can be both spontaneous and stimulated Raman scattering processes, but, for our channel distribution and power levels, we just need to worry about spontaneous Raman scattering. Spontaneous Raman scattering, in turn, can present Stokes and anti-Stokes processes. In Stokes processes a photon produces a lower energy photon and a phonon. In anti-Stokes processes, a photon and an existing phonon in the fiber produce a new photon with a higher energy (a lower wavelength). The pump laser signal and the scattered photon can go in the same direction (copropagation) or in opposite directions (counterpropagation). We are considering unidirectional fiber links, where duplex communication is done with one fiber for each direction, and with a pump, the classical channel, at a much higher wavelength than the quantum channels. In that scenario, the most important photon contamination in the quantum channel are the photons coming from anti-Stokes spontaneous Raman scattering of the photons of the classical channel copropagating with the quantum channel. For our frequencies and channel powers, the Raman noise power can be approximated as [9, 10]:
$\text{RNP}=P_{0}\frac{\beta_{q}}{\alpha_{q}-\alpha_{c}}\left(e^{-\alpha_{c}z}-e^{-\alpha_{q}z}\right),$ (1)
[…]
In the standard silica fiber we consider, we have an attenuation $\alpha_{cdB}=$0.17\text{\,}\mathrm{d}\mathrm{B}\mathrm{/}\mathrm{k}\mathrm{m}$\$, corresponding to $\alpha_{c}=$0.0391\text{\,}\mathrm{k}\mathrm{m}^{-1}$\$ for the classical channels around $1550\text{\,}\mathrm{nm}$ and $\alpha_{qdB}=$0.32\text{\,}\mathrm{d}\mathrm{B}\mathrm{/}\mathrm{k}\mathrm{m}$\$, $\alpha_{q}=$0.0737\text{\,}\mathrm{k}\mathrm{m}^{-1}$\$, for the quantum channels around $1310\text{\,}\mathrm{nm}$.
In Equation (1), we can rewrite the term giving the evolution with distance $e^{-\alpha_{c}z}-e^{-\alpha_{q}z}$ as $e^{-\alpha_{c}z}\left(1-e^{-(\alpha_{q}-\alpha_{c})z}\right)$. For our attenuation constants and total distances above $50\text{\,}\mathrm{km}$, the second exponential introduces a correction in the Raman power below the 20%. Depending on the scenario, we could, for simplicity, drop the second exponential in the model to give an upper bound for the Raman noise, but the term can be included if there are many short links. This approximation could be useful if we wanted to collect multiple nonlinear effects in one term. We could then plug all the terms into one constant and reduce the computation complexity when approximated linearizations give a good enought estimation. As the computational cost is small for the Raman formula we have discussed, we will keep the full term.
For multiple channels we can just sum the contribution of each classical pump field [11] assuming that we have a constant attenuation and spontaneous Raman scattering coefficient for all the signals in the C band. In the final term, nonlinear contamination for each different optical fiber can be reduced to a constant $\gamma_{nl}$ multiplied by the input power and the attenuation exponentials. If needed, the model can also be easily adapted to other scenarios by adding a term for the counterpropagating signals or distinct values of $\gamma_{nl}$ for heterogeneous networks with fibers with different nonlinear responses.
3.2 Classical and quantum signal to noise ratios
@@
We will use the signal-to-noise ratio of the classical and quantum channels as our guiding quality metric.
@@
In the classical channels we use the usual definition for the signal-to-noise ratio $SNR=P/N$ as the ratio between the signal power $P$ and the noise power $N$. For values below a certain threshold, which depends on the receivers and modulation schemes, we assume the signal cannot be interpreted correctly. If we have a fixed power at each transmitter, the receivers at the end of shorter links will receive a higher power. We are going to assume there is a fixed noise power for each link, mostly related to thermal noise at the receivers. For each channel we have
$\text{SNR}=\frac{P}{N}=\frac{e^{-\alpha_{c}L_{c}}P_{tx,c}}{N},$ (2)
where $\alpha_{c}$ is the attenuation at the transmission wavelength in natural units ( $\text{\,}\mathrm{k}\mathrm{m}^{-1}$), $L_{c}$ is the total length of the path we use for the channel, $P_{tx,c}$ is the launch power of the classical signal and $N$ is the noise term. All the signal and noise powers are normalized to a reference level of $1\text{\,}\mathrm{mW}$ in the simulations.
@@
We estimate the effects of the noise on the quantum channel by defining a quantum signal-to-noise ratio that compares the power of the quantum channel at the receiver and all the noise during detection, including the photons coming from nonlinear effects. We define the formula
$\text{QSNR}=\frac{e^{-\alpha_{q}L_{q}}P_{tx,q}}{N_{F}+\gamma_{nl}\sum_{i}P_{tx,i}\cdot\left(e^{-\alpha_{c}L_{i}}-e^{-\alpha_{q}L_{i}}\right)}.$ (3)
Here, $\alpha_{q}$ is the attenuation of the fiber at the transmission wavelength of the quantum signal (in  $\text{\,}\mathrm{k}\mathrm{m}^{-1}$), $L_{q}$ is the length of the route allocated to quantum channel $q$, $P_{tx,q}$ is the transmitted power of that quantum channel, $N_{F}$ is a fixed noise term, $\gamma_{nl}$ is a fixed nonlinear coefficient term that can be adapted to the physical parameters of the fiber used in our particular network, $P_{tx,i}$ represents the classical signal launch power at the beginning of each shared span $i$ with shared length $L_{i}$ and $\alpha_{c}$ represents the attenuation coefficient at the transmission wavelength of the classical channel in  $\text{\,}\mathrm{Np}\mathrm{/}\mathrm{k}\mathrm{m}$. The shared span is given by $L_{i}$, which corresponds to the length of the shared fiber. The sum includes all the fiber links in the path of the quantum signal in which there are classical channels. We will add one term to the sum for each classical channel sharing that link.
We consider all the powers in the QSNR are normalized to the single photon level for the corresponding wavelength in the O-band so that we can interpret them as a photon number. The numerator in Equation (3) gives an estimation of the expected photon number that reaches the receiver in a detection window. It is computed from the transmitted photon number and the attenuation in the journey to the destination node. We will also include in that term effects that are equivalent to losses, like detection efficiencies below 1, which are common for the simplest single photon detectors.
Single photon detectors are a key element in discrete variable QKD. The two most important parameters are their quantum efficiency $\eta$ and their dark count rate, DCR. The detector efficiency gives the fraction of photons that reach the photon detector that are actually converted into a macroscopical electrical pulse that can be detected. Dark counts give the rate at which the detector produces a false count (a phonon or any thermal process triggers the same response as a transmitted photon would when there are none). At the usual telecom wavelengths, most affordable single photon detectors have a low detection efficiency, especially if we want to keep dark counts at bay [31].
Many systems use avalanche photodiodes which are notoriously inefficient at the IR telecom bands, where 10-20% efficiencies are the norm. $P_{tx,q}$ will be set to $\mu\eta$, where $\eta$ is the quantum efficiency of the detector and $\mu$ is the average photon number at the output of a transmitter generating weak coherent states with an amplitude giving a mean photon number $\mu<1$. The value of $\mu$ can be adapted to achieve the optimal key rate for a particular protocol or can be given a fixed value, typically much smaller than one to avoid photon number splitting attacks.
@@
The fixed amount $N_{F}$ that collects thermal photons and equivalent effects, like dark counts in the detectors (false detections caused by phonons or other effects that are not related to the incoming light). In our experiments we consider the dominant fixed source of noise are dark counts, which happen in each detection bin with a probability $p_{d}$.
We estimate $N_{c}$ with the simplified model we have described in Section 3 and write the total noise as the product of a non-linear coefficient $\gamma_{nl}$ by the sum of the classical power of each classical channel multiplied by its corresponding exponential distance terms for the shared total distance with the quantum channel of interest.
In the general model, $\gamma_{nl}$ is left as an adjustable parameter and, for any physical QKD network, its value will depend on the detectors we use, the physical properties of the optical fiber in our network and the parameters of the QKD protocol such as time widths for detection or the transmitted number of photons.
We will give an example system as a reference to show how to set the parameters for a concrete scenario and study the changes in the achievable key rate for a specific QKD protocol, BB84. The parameters can also be studied independently if the focus is on the effect of different network topologies or more general scenarios.
In order to see how many photons are scattered into the O-band and detected at the QKD receiver we can go to Equation (1). The number of photons that produce a false count are
$\eta\text{RNP}\frac{\Delta T}{h\nu}\approx\eta\frac{\beta_{q}}{\alpha_{q}-\alpha_{c}}\left(e^{-\alpha_{c}z}-e^{-\alpha_{q}z}\right)\frac{\Delta T}{h\nu}P_{0}=\gamma_{nl}P_{0}\left(e^{-\alpha_{c}z}-e^{-\alpha_{q}z}\right)$ (4)
where $\Delta T$ is the time of the detection window at the receiver and $\eta$ is the quantum efficiency of the detector and we divide the input noise power by the energy of a single photon at $1550\text{\,}\mathrm{nm}$ to get the noise photon rate reaching the detector.
We take a reference QKD system with a detector that, for a quantum efficiency $\eta=0.1$, has a dark count probability $p_{d}=10^{-4}$ per detection bin, corresponding to a time window $\Delta T=$10\text{\,}\mathrm{ns}$\$ and a detector with a dark count rate of $10\text{\,}\mathrm{kHz}$. We normalize input powers so that $P_{0}=1$ for a $1\text{\,}\mathrm{mW}$ classical signal at $1550\text{\,}\mathrm{nm}$ and absorb the $7.8\cdot 10^{7}$ photons per time bin of such a signal into the constant $\gamma_{nl}$. We consider an optical fiber with a constant Raman coefficient for the portion of the O band we measure.
For all these parameters and an optical fiber with a Raman coefficient $\beta_{q}=\beta\delta\lambda=${10}^{-12}\text{\,}\mathrm{k}\mathrm{m}^{-1}$\$ at $1310\text{\,}\mathrm{nm}$, consistent with measured values for typical fibers with a filter of $100\text{\,}\mathrm{GHz}$ [32], we get $\gamma_{nl}=2.25\cdot 10^{-4}$.
@@
This approximation gives a reasonable estimation of which classical channels are likely to introduce the most noise without a detailed physical model of each situation. The parameters can be adjusted to model each particular scenario, including using different non-linear coefficients for links with different optical fibers or when the classical and quantum signals are copropagating or counterpropagating.
3.3 Quantum key rate example
QKD exchanges start with an initial phase where the quantum states are sent through the optical fiber and measured. The results of the measurements are used to estimate the QBER for the current state of the fiber link and the devices at the receiver. After this initial estimation, there is a phase of error correction and privacy amplification where the bits from the raw key, which are directly related to the measurement results, are modified to produce a smaller secure key which should be the same for both parties. The size of this final key depends on the estimated information a potential eavesdropper can have, assuming any imperfection during the quantum state transmission and measurement has been caused by an active adversary.
The rate at which we can produce a secure key depends on the protocol, the link distance and certain parameters of the channel and the equipment. The attenuation in the optical fiber $\alpha$ and the efficiency $\eta$ and dark count rate $p_{d}$ of the detector are usually the most important variables.
Each QKD protocol has different security proofs that bound what an adversary can do under different assumptions. The best models that give tighter bounds and avoid throwing away more bits than necessary can be quite involved and depend on the particular QKD protocol [33]. There are also different numerical techniques that optimize the way the key is extracted from the measurements [34]. While we give a general description in terms of the QSNR, which gives a good intuition of the expected improvement for any protocol, in this Section we give a simple example of key rate values using the simplified key rate formula for the BB84 protocol under collective combined photon-number-splitting and cloning attacks as described in [35]. Under this scenario, we can approximate the maximum achievable secure key rate as
$R\!=\!\frac{1}{2}\!\left(\mu t\eta+2p_{d}\right)\!\left[1\!-\!H(\text{QBER})\right]-\frac{1}{2}\mu t\eta\left[\!\left(t-\frac{\mu}{2}\right)\!I_{1}(D_{1})+\frac{\mu}{2}\right]$ (5)
for a quantum bit error rate (QBER)
$\text{QBER}=\frac{1}{2}-\frac{V}{2\left(1+\frac{2p_{d}}{\mu t\eta}\right)},$ (6)
for a system using a detector with efficiency $\eta$ and a dark count probability per gate $p_{d}$. Here,
$I_{1}(D_{1})=1-H\left(\frac{1}{2}+\sqrt{D_{1}(1-D_{1})}\right)$
with $D_{1}=\frac{1-V}{2-\mu/t}$, where $V$ is the visibility in the interferometric detector, which can be established experimentally during the protocol setup, and
$H(\text{QBER})=-\log_{2}(\text{QBER})\text{QBER}-\log_{2}(1-\text{QBER})(1-\text{QBER})$
is the binary Shannon entropy for the QBER.
In our model, we collect the noise photons coming from nonlinear mixing of the classical signals into an effective dark count rate
$p_{d}^{\prime}=N_{F}+\gamma_{nl}\sum_{i}P_{tx,i}\cdot\left(e^{-\alpha_{c}L_{i}}-e^{-\alpha_{q}L_{i}}\right),$ (7)
that will include a fixed number of photons from dark counts $N_{F}=p_{d}$ and the photons produced by nonlinear processes for each possible route between origin and destination. The constant $\gamma_{nl}$ absorbs all the fiber parameters and the detector efficiency $\eta$ so that we can get the expected number of noise photon detections in a single detection time bin.
This corrected formula gives the number of “false” detections that are not triggered by a photon from Alice. As in other QKD security proofs, the rate formula we use assumes all the imperfections are caused by the presence of an eavesdropper.
All the other parameters being equal, the QSNR decreases as the signal photons in the quantum channel attenuate during propagation. As soon as the losses and other imperfections in the channel reach a critical level, we can no longer guarantee we can extract a private key that cannot be guessed, totally or partially, by an attacker. This gives a distance limit for secure key generation.
4 Network Model